Bag C and Bag D each contain 70 marbles. All the marbles inside both bags are red, white or blue.

In bag C, R:W = 2:3 and W:B = 3:5
In bag D, R:W = 2:3 and W:B = 4:5

What is the total number of white marbles in BOTH bags?
a. 13
b. 21
c. 24
d. 45

Answer by jim_thompson5910(35256)

Variables

Rc = number of red marbles in bag C
Wc = number of white marbles in bag C
Bc = number of blue marbles in bag C

Rd = number of red marbles in bag D
Wd = number of white marbles in bag D
Bd = number of blue marbles in bag D


Bag C:
Ratio 1
R:W = 2:3
Rc:Wc = 2:3
Rc/Wc = 2/3
3Rc = 2Wc
2Wc = 3Rc
Wc = 3Rc/2


Bag C:
Ratio 2
W:B = 3:5
Wc:Bc = 3:5
Wc/Bc = 3/5
5Wc = 3Bc
5(3Rc/2) = 3Bc
15Rc/2 = 3Bc
15Rc = 2*3Bc
15Rc = 6Bc
6Bc = 15Rc
Bc = 15Rc/6


Bag C and Bag D each contain 70 marbles
Bag C contains 70 marbles
(number of red marbles)+(number of white marbles)+(number of blue marbles) = total
(Rc) + (Wc) + (Bc) = 70
(Rc) + (3Rc/2) + (15Rc/6) = 70
6 = 6*70
6*(Rc) + 6*(3Rc/2) + 6*(15Rc/6) = 6*70
6Rc + 9Rc + 15Rc = 420
30Rc = 420
Rc = 420/30
Rc = 14
There are 14 red marbles in bag C
Wc = 3Rc/2
Wc = 3*14/2
Wc = 21
There are 21 white marbles in bag C

=============================================================================================
Bag D
In bag D, R:W = 2:3 and W:B = 4:5


Bag D
Ratio 1
R:W = 2:3
Rd:Wd = 2:3
Rd/Wd = 2/3
3Rd = 2Wd
Wd = 3Rd/2

Bag D
Ratio 2
W:B = 4:5
Wd:Bd = 4:5
Wd/Bd = 4/5
5Wd = 4Bd
5(3Rd/2) = 4Bd
15Rd/2 = 4Bd
Bd = 15Rd/8

Bag D also contains 70 marbles
(number of red marbles)+(number of white marbles)+(number of blue marbles) = total
(Rd) + (Wd) + (Bd) = 70
(Rd) + (3Rd/2) + (15Rd/8) = 70
8 = 8*70
8Rd + 12Rd + 15Rd = 560
35Rd = 560
Rd = 560/35
Rd = 16
There are 16 red marbles in bag D
Wd = 3Rd/2
Wd = 3*16/2
Wd = 24
There are 24 white marbles in bag D


Wc + Wd = 21 + 24 = 45
So there are 45 white marbles total.